OpenStudy (bbb911):

This is Geometry: Given that two sides of a triangle are 20 and 66. Find the range of possible measures for the third side, s.

OpenStudy (anonymous):

i don't have enough yarn for this.

OpenStudy (anonymous):

s>9000

OpenStudy (anonymous):

okay someone gave me a medal for being funny, so i'll give this question my best shot

OpenStudy (anonymous):

that was me

OpenStudy (bbb911):

first answer sooo wrong sorry i have multiple choice

OpenStudy (anonymous):

i was joking sorry

OpenStudy (anonymous):

what are the choices i can tell you which one it is

OpenStudy (anonymous):

46<s<86

OpenStudy (anonymous):

20^2+66^2=c^2 or 20^2+b^2=66^2

OpenStudy (anonymous):

i think, geometry was awhile ago

OpenStudy (anonymous):

i just have no clue how to solve this algebraically outside of my head

OpenStudy (bbb911):

yay thanks akbr0

OpenStudy (anonymous):

OH DUHHH A2+B2=C2

OpenStudy (anonymous):

TIME FOR A REAL SHOWER. I THINK I'LL HAVE MY WHISKEY NOW

OpenStudy (anonymous):

(btw, my 2's up there are squares)

OpenStudy (anonymous):

if 66 is the longest side then the sum of the other two sides must be > 66. s+20>66 s>46 if s is the longest side then the sum of the other two sides are greater than s: s<20+66 s<86 46<s<86

OpenStudy (bbb911):

thank you too rsvital

OpenStudy (anonymous):

youre welcome